Calculate the osmolality for each of these six solutions 0.1, 0.3, 0.5, 0.65, 0.85, 1% NaCl. Which concentration is closest to that of plasma (290 mOsm/kg)?
1
Expert's answer
2013-09-23T10:49:14-0400
Osmolarity: 290 mOsm/kg.
MW of NaCl: 58.5
Concentration: 0.1% solution.
Calculations:
0.1g/100ml or 1 mg/ml. 1/58.5 = 0.017 x valence(1) = 0.017 meq/ml. Osmolarity: 1 g/Liter =0.017 moles or 17 mmol x species(2)= 34 mOsm/liter
Concentration: 0.3% solution.
Calculations:
0.3g/100ml or 3mg/ml. 3/58.5 = 0.051 x valence(1) = 0.051 meq/ml. Osmolarity: 3 g/Liter =0.051 moles or 51 mmol x species(2)= 102 mOsm/liter
Concentration: 0.5% solution.
Calculations:
0.5g/100ml or 5 mg/ml. 5/58.5 = 0.085 x valence(1) = 0.085 meq/ml. Osmolarity: 5 g/Liter =0.085 moles or 85 mmol x species(2)= 170 mOsm/liter
Concentration: 0.65% solution.
Calculations:
0.65g/100ml or 6,5 mg/ml. 6,5/58.5 = 0.11 x valence(1) = 0.11 meq/ml. Osmolarity: 6,5g/Liter =0.11 moles or 110 mmol x species(2)= 220 mOsm/liter
Concentration: 0.85% solution.
Calculations:
0.85g/100ml or 8,5 mg/ml. 8,5/58.5 = 0.145 x valence(1) = 0.145 meq/ml. Osmolarity: 8,5 g/Liter =0.145 moles or 145 mmol x species(2)= 290 mOsm/liter
Concentration: 1% solution.
Calculations:
1g/100ml or 10mg/ml. 10/58.5 = 0.17 x valence(1) = 0.17 meq/ml. Osmolarity: 10 g/Liter =0.17 moles or 170 mmol x species(2)= 340 mOsm/liter
Answer: concentration of solution of NaCl = 0,85 is closest to that of plasma